Equipment and Calibration

Equipment history at East End

| ALL | | Anemometer || Barometer || Battery Bank || Camera || CO2 Probe || Conductivity Sensor || Data Logger || DO Sensor || Gas Analyzer || Heat Flux Plate || Interface || Multiplexer || Narrow Band || Net Radiometer || PAR Sensor || Precipitation || RH & Temp || Solar Controller || Solar Panel || Spectrometer || Water Level |

SELECT a.recno, a.eh_date, a.eh_loc1, a.eh_loc2, IF(a.eh_loc2=:loc, a.eh_locref2, a.eh_locref1) AS ref, a.eh_locref2, a.eh_title, a.eh_notes, a.eh_id, b.serialno, b.equip, b.modelno FROM equiphist a INNER JOIN equiplist b ON a.eh_id=b.recno WHERE (a.eh_loc1=:loc1 OR a.eh_loc2=:loc2) AND b.equip=:eq ORDER BY b.serialno, a.eh_date

Date Serial No. Model Action Ref Notes
RH & Temp
2024-08-29 E5905 HygroVUE10 Installed RH & Temp Installed at EE to replace the HMP60 sn L1740420 that was finally starting to die
2013-11-25 G4540023 HMP60 Moved RH & Temp The DWR Corn / Construction tower was decommissioned and rechristened at its new permanent location at the Twitchell East End Wetland
2015-01-05 G4540023 HMP60 Removed RH & Temp Removed from the site due to the tower falling over in the wind on 2014-12-30. Likely undamaged.
2015-01-20 G4540023 HMP60 Installed RH & Temp Reinstalled after being removed when the scaffolding fell over.
2017-12-18 G4540023 HMP60 Removed RH & Temp Removed from Eat End for a 24hour side-by-side comparison at Twitchell Alfalfa
2017-12-19 G4540023 HMP60 Removed RH & Temp Removed from the HMP inter-comparison at Twitchell Alfalfa and returned to East End
2019-01-24 G4540023 HMP60 Removed RH & Temp Brought back to the lab from EE for inspection b/c of data drops/spikes. It was replaced by HMP45 sn U3030046.
2019-01-31 L1740420 HMP60 Installed RH & Temp Installed at EE to replace HMP45 sn U3030046, which was removed for bad RH readings.
2024-08-29 L1740420 HMP60 Removed RH & Temp Removed from EE and brought back to the lab b/c the sensor is starting to die. It was replaced by Hygrovue sn E5905.
2019-01-24 U3030046 HMP45AC Installed RH & Temp Installed at EE to replace G4540023, which was brought back to the lab for inspection.
2019-01-31 U3030046 HMP45AC Removed RH & Temp Removed from EE b/c RH was not working (ranged only b/t 10 and 13%). It was replaced by HMP60 sn L1740420.
Export to BADM